# Model-free policy gradient for discrete-time mean-field control

### v1

@@ 3.4 Error of the Estimator
$\left(\Sigma_{t}^{\theta,\varepsilon}\right)_{i,j}\coloneqq{\mathrm{Var}}\left[\left(\widehat{\nabla_{\theta}l_{t}^{\theta}}\right)_{i,j}\right],\quad i=1,\ldots,d,j=1,\ldots,D.$
@@
Let Assumptions 1, 2 and 3 hold. Let $\varepsilon\in(0,1)$, $\theta\in\Theta$, $n\in\mathbb{N}$ and $N\in\mathbb{N}$. Let $\widehat{\nabla V}(\theta)^{(1)},\ldots,\widehat{\nabla V}(\theta)^{(N)}$ be NN independent copies of the policy gradient estimator defined in (3.6), where each estimator is computed using an independent version of the estimators $\left(\widehat{\nabla_{\theta}l_{t}^{\theta}}\right)_{t=0}^{T}$ defined in (3.4). and let $\widehat{\nabla V}(\theta)=N^{-1}\sum_{k=1}^{N}\widehat{\nabla V}(\theta)^{(k)}$ be the averaged estimator. Then, for all $\theta\in\Theta$,
$\mathrm{MSE}(\theta,N,n,\varepsilon)\coloneqq\max_{j=1,\ldots,D}\mathbb{E}\left[\left|\widehat{\nabla V}(\theta)_{j}-\nabla_{\theta}V(\mu,\theta)_{j}\right|^{2}\right]\leq\mathcal{K}_{1,\varepsilon}\varepsilon^{2}+\frac{\mathcal{K}_{2,\varepsilon}}{N\varepsilon^{2}}+\frac{\mathcal{K}_{3,\varepsilon}}{nN\varepsilon^{4}}\sum_{t=0}^{T}\mathcal{C}_{t}\varepsilon^{-2t},$ (3.9)
@@
$\mathcal{K}_{2,\varepsilon}$ $=2(T+1)^{3}M_{0}^{2}\sum_{t=0}^{T}\left(E_{t}(\theta,\varepsilon)+\left\|\nabla_{\theta}l_{t}^{\theta}\right\|+\varepsilon C\right)$
@@ 4.1 Two-State Two-Action Toy Problem
where $W_{1}$ is the 1-Wasserstein distance, $\lambda>0$ is a fixed scalar parameter and $B$ is a binomial distribution with parameter $p$ such that $1-\lambda_{0}\leq p\leq\lambda_{1}$. The terminal reward function is $g(x,\mu)=r(x,\mu)$. Although the original problem in [17] is an infinite-horizon discounted problem, we consider here a finite-horizon version of with time horizon $T=2$. In this setting, there exists an optimal stationary policy given by
@@
In our experiments, we take the following parameters: $\lambda_{0}=0.5,\lambda_{1}=0.8,\lambda=10,p=0.6$. This policy guarantees that, no matter the initial distribution $\mu_{0}$, the distribution at all time steps $t\geq 1$ is $\mu_{t}^{\star}=B$. The training and evaluation are done as follows. For each training episode we simulate the population starting from a random initial distribution $\mu\in\mathcal{P}(\mathcal{X})^{*}$ by drawing $\mu(L)\sim\mathcal{U}([0.1,0.9])$. Every 10 training episodes, we freeze the policy and sample a validation episode for which we compute the population reward $V(\mu_{0})$ starting from a fixed initial distribution $\mu_{0}=(\mu_{0}(0),\mu_{0}(1))=(0.2,0.8)$. During training, we compute $N=200$ trajectories for MF-REINFORCE, with $n=10$ trajectories for the gradient of logits estimation. For this first example, the policy considered is a simple static policy that outputs a $|\mathcal{X}|\times|\mathcal{A}|$ matrix corresponding to the probability of each action given each state. We make this particular choice of policy to check whether MF-REINFORCE is capable of recovering the optimal static policy given in (4.1) We use the Adam optimizer [24] and train for 5,000 episodes using a learning rate of $10^{-3}$. The experiments are run for $\varepsilon$ ranging in the set $\{0.2,0.5,1.0,2.0\}$. The validation rewards for the different values of $\varepsilon$ are shown in Figure 1. One can see that a larger value of $\varepsilon$ leads to faster increase in the value function at first, but yields worse final policies. In particular, we see that choosing $\varepsilon$ too large can lead to catastophic failure. To evaluate how well the learned policies match the optimal static policy $\pi^{\star}$ (4.1), we compute the average absolute errors in the resulting estimates of $\pi(\mathrm{ST}\mid 0)$ and $\pi(\mathrm{ST}\mid 0)$ over 5 independent training runs for each value of $\varepsilon$ considered. The results are reported in Table 1. These results match adequately with the curves of Figure 1: the smaller $\varepsilon$, the better and closer to optimal the learned policy is.
@@ 4.3 Distribution Planning Example
Finally, we consider a problem similar to the distribution planning example of [6], Section 6.2. The goal is to match a fixed target distribution $\mu_{\mathrm{target}}\in\mathcal{P}(\mathcal{X})$, where $\mathcal{X}=\mathbb{Z}/10\mathbb{Z}$. The action space is $\mathcal{A}=\{-1,0,+1\}$, and the transition probability is deterministic $P(\cdot\mid x,a,\mu)=\delta_{x+a}$. At each step, we penalize the deviation of the population from the target distribution and penalize when an agent chooses to move. Specifically, the step reward is $r(x,a,\mu)=-\lambda|a|-\sum_{x\in\mathcal{X}}|\mu(x)-\mu_{\mathrm{target}}(x)|^{2}$, where $\lambda>0$, and the terminal reward is $g(x,\mu)=-\sum_{x\in\mathcal{X}}|\mu(x)-\mu_{\mathrm{target}}(x)|^{2}$. In this example, we take the length of episodes to be $T=5$ (for both training and validation) and $\lambda=0.01$. For each training episode, we simulate the population starting from a random initial distribution $\mu\in\mathcal{P}(\mathcal{X})^{*}$. Every 10 training episodes, we freeze the policy and sample a validation episode, for which we compute the population reward $V(\mu_{0})$ starting from a fixed initial distribution $\mu_{0}=\mathcal{U}(\mathcal{X})$. During training, we compute $N=500$ trajectories for MF-REINFORCE, with $n=10$ trajectory for the gradient of logits estimation. The policy is given by a 2-layer MLP, with 256 hidden units and $\tanh$ activation. We use the Adam optimizer and train for 100,000 episodes using a learning rate of $10^{-4}$. The experiments are run for $\varepsilon$ ranging in the set $\{0.5,0.75,1.0,2.0\}$. In this example, we see that taking a large value of $\varepsilon$, i.e., $\varepsilon\in\{1,2\}$, leads to more stable training and faster convergence, as shown in Figure 4. However, the resulting trajectories still differ qualitatively between $\varepsilon=1.0$ and $\varepsilon=2.0$, as seen in Figure 5.
@@ 5.1 Proofs of Results of Section 2
$\nabla_{\theta}V(\mu,\theta)=\nabla_{\theta}\int\left(\sum_{t=0}^{T-1}r(x_{t},a_{t},\mu_{t}^{\theta})+g(x_{T},\mu_{T}^{\theta})\right)P_{\theta}({\mathrm{d}}x_{0},da_{0},\ldots,{\mathrm{d}}a_{T-1},{\mathrm{d}}x_{T})$ (5.1)
@@
$+\int\left(\sum_{t=0}^{T-1}r(x_{t},a_{t},\mu_{t}^{\theta})+g(x_{T},\mu_{T}^{\theta})\right)\nabla_{\theta}P_{\theta}({\mathrm{d}}x_{0},da_{0},\ldots,{\mathrm{d}}a_{T-1},{\mathrm{d}}x_{T}).$
@@
$=\int\left(\sum_{t^{\prime}=0}^{T-1}\nabla_{\theta}\log p(\theta,a_{t^{\prime}},t^{\prime},x_{t^{\prime}},\mu_{t^{\prime}}^{\theta})+\left(\nabla_{l}\log\tilde{p}(\theta,a_{t^{\prime}},t^{\prime},x_{t^{\prime}},l_{t^{\prime}}^{\theta})\tilde{P}(x_{t^{\prime}+1}\mid x_{t^{\prime}},a_{t^{\prime}},l_{t^{\prime}}^{\theta})\right)\nabla_{\theta}l_{t^{\prime}}^{\theta}\right)$
$\cdot\left(\sum_{t=0}^{T-1}r(x_{t},a_{t},\mu_{t}^{\theta})+g(x_{T},\mu_{T}^{\theta})\right)P_{\theta}({\mathrm{d}}x_{0},{\mathrm{d}}a_{0},\ldots,{\mathrm{d}}a_{T-1},{\mathrm{d}}x_{T}),$
@@
$\mathrm{RF}(\theta)\coloneqq\int\left(\sum_{t^{\prime}=0}^{T-1}\nabla_{\theta}\log p(\theta,a_{t^{\prime}},t^{\prime},x_{t^{\prime}},\mu_{t^{\prime}}^{\theta})\right)\left(\sum_{t=0}^{T-1}r(x_{t},a_{t},\mu_{t}^{\theta})+g(x_{T},\mu_{T}^{\theta})\right){\mathrm{d}}P_{\theta},$ (5.2)
@@
By performing the change of variables $\tilde{\lambda}_{t}=l^{\pi}_{t}+\varepsilon\lambda_{t}$, we have
@@
$\cdot\mu_{0}({\mathrm{d}}y_{0})\prod_{\tau=0}^{T-1}\mathrm{Leb}({\mathrm{d}}\tilde{\lambda}_{\tau})\tilde{\pi}({\mathrm{d}}a_{\tau}\mid\tau,y_{\tau},\tilde{\lambda}_{\tau})\tilde{P}({\mathrm{d}}y_{\tau+1}\mid y_{\tau},a_{\tau},\tilde{\lambda}_{\tau}).$
Now, let’s consider a parametrized family of policies $\pi_{\theta}$ with densities $p_{\theta}$, for $\theta\in{\mathbb{R}}^{D}$. We write $V_{\varepsilon}(\mu,\theta)=V_{\varepsilon}^{\pi_{\theta}}(\mu)$, $\mu_{t}^{\theta}=\mu_{t}^{\pi_{\theta}}$, $l_{t}^{\theta}=l_{t}^{\pi_{\theta}}$, $\alpha_{t}^{\theta,\varepsilon}=\alpha_{t}^{\pi_{\theta},\Lambda,\varepsilon}$ $Y_{t}^{\theta,\varepsilon}=Y_{t}^{\pi_{\theta},\Lambda,\varepsilon}$. The differentiability of $V_{\varepsilon}(\mu,\theta)$ follows from the differentiability of $\tilde{p}$ in $\theta$ and $l$ and the differentiability of $\tilde{P}$ in $l$. Then, given the assumptions on $r,g,\pi_{\theta}$, one can differentiate under the integral sign to get
@@
$\cdot\sum_{t=0}^{T}\varepsilon^{-2}(l_{t}^{\theta}-\tilde{\lambda}_{t})\nabla_{\theta}l_{t}^{\theta}\prod_{\tau=0}^{T}\frac{1}{(2\pi)^{(d-1)/2}\varepsilon^{d-1}}\exp\left(-\frac{\|\tilde{\lambda}_{\tau}-l_{\tau}^{\theta}\|^{2}}{2\varepsilon^{2}}\right)$
$\cdot\mu_{0}({\mathrm{d}}y_{0})\prod_{\tau=0}^{T-1}\mathrm{Leb}({\mathrm{d}}\tilde{\lambda}_{\tau})\tilde{\pi}_{\theta}({\mathrm{d}}a_{\tau}\mid\tau,y_{\tau},\tilde{\lambda}_{\tau})\tilde{P}({\mathrm{d}}y_{\tau+1}\mid y_{\tau},a_{\tau},\tilde{\lambda}_{\tau})$
$+\int\left(\sum_{t=0}^{T-1}\tilde{r}(y_{t},a_{t},\tilde{\lambda}_{t})+\tilde{g}(y_{T},\tilde{\lambda}_{T})\right)\prod_{\tau=0}^{T}\frac{1}{(2\pi)^{d/2}\varepsilon^{d}}\exp\left(-\frac{\|\tilde{\lambda}_{\tau}-\mu_{\tau}^{\theta}\|^{2}}{2\varepsilon^{2}}\right)$
$\cdot\nabla_{\theta}\left(\mu_{0}({\mathrm{d}}y_{0})\prod_{\tau=0}^{T-1}\mathrm{Leb}({\mathrm{d}}\tilde{\lambda}_{\tau})\tilde{\pi}_{\theta}({\mathrm{d}}a_{\tau}\mid\tau,y_{\tau},\tilde{\lambda}_{\tau})\tilde{P}({\mathrm{d}}y_{\tau+1}\mid y_{\tau},a_{\tau},\tilde{\lambda}_{\tau})\right).$
@@
$=\mu({\mathrm{d}}y_{0})\sum_{t=0}^{T-1}\nabla_{\theta}\log\tilde{p}(\theta,a_{t},t,y_{t},\tilde{\lambda}_{t})\prod_{\tau=0}^{T-1}\mathrm{Leb}({\mathrm{d}}\tilde{\lambda}_{\tau})\tilde{\pi}_{\theta}({\mathrm{d}}a_{\tau}\mid\tau,y_{\tau},\tilde{\lambda}_{\tau})\tilde{P}({\mathrm{d}}y_{\tau+1}\mid y_{\tau},a_{\tau},\tilde{\lambda}_{\tau}).$
@@
$\cdot\left(\sum_{t=0}^{T}\varepsilon^{-1}\Lambda_{t}\left(\nabla_{\theta}l_{t}^{\theta}\right)\right)\Bigg]$
$+\mathbb{E}\Bigg[\left(\sum_{t=0}^{T-1}\tilde{r}(Y^{\theta,\varepsilon}_{t},\alpha_{t}^{\theta,\varepsilon},l_{t}^{\theta}+\varepsilon\Lambda_{t})+\tilde{g}(Y_{T}^{\theta,\varepsilon},l_{T}^{\theta}+\varepsilon\Lambda_{T})\right)$
$\cdot\left(\sum_{t=0}^{T-1}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta,\varepsilon},Y_{t}^{\theta,\varepsilon},l_{t}^{\theta}+\varepsilon\Lambda_{t}\right)\Bigg].$
@@
Let’s decompose the expression (2.5) of the gradient of the perturbed value function,
@@
$\to\mathbb{E}\Bigg[\left(\sum_{t=0}^{T-1}r(X^{\theta}_{t},\alpha_{t}^{\theta},\mathbb{P}_{X_{t}^{\theta}})+g(X_{T}^{\theta,\varepsilon},\mathbb{P}_{X^{\theta}_{T}})\right)\left(\sum_{t=0}^{T-1}\nabla_{\theta}\log p(\theta,\alpha_{t}^{\theta},t,X_{t}^{\theta},\mathbb{P}_{X^{\theta}_{t}})\right)\Bigg]=\mathrm{RF}(\theta).$
We now look at the first expectation in (5.3) as $\varepsilon\to 0$.
@@
$M_{\varepsilon}(\theta)$ $\coloneqq\mathbb{E}\left[\left(\sum_{t=0}^{T-1}\tilde{r}(Y^{\theta,\varepsilon}_{t},\alpha_{t}^{\theta,\varepsilon},l_{t}^{\theta}+\varepsilon\Lambda_{t})+\tilde{g}(Y_{T}^{\theta,\varepsilon},l_{T}^{\theta}+\varepsilon\Lambda_{T})\right)\cdot\left(\sum_{t=0}^{T}\varepsilon^{-1}\Lambda_{t}\nabla_{\theta}l_{t}^{\theta}\right)\right]$
$=\mathbb{E}\left[\left(\sum_{t=0}^{T-1}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta}+\varepsilon\Lambda_{t})+\tilde{g}(X_{T}^{\theta},l^{\theta}_{T}+\varepsilon\Lambda_{T})\right)\cdot\left(\sum_{t=0}^{T}\varepsilon^{-1}\Lambda_{t}\nabla_{\theta}l_{t}^{\theta}\right)R_{T}^{\theta,\varepsilon}(\theta)\right],$
@@
$\cdot\left(1+\varepsilon\nabla_{l}\log\tilde{P}(X_{\tau+1}^{\theta}\mid X_{\tau}^{\theta},\alpha_{\tau}^{\theta},l_{\tau}^{\theta})\cdot\Lambda_{\tau}+o(\varepsilon)\right)\Big]$

### v2

@@ 3.4 Error of the Estimator
$\left(\Sigma_{t}^{\theta,\varepsilon}\right)_{i,j}\coloneq{\mathrm{Var}}\left[\left(\widehat{\nabla_{\theta}l_{t}^{\theta}}\right)_{i,j}\right],\quad i=1,\ldots,d,j=1,\ldots,D.$
@@
Let Assumptions 1, 2 and 3 hold. Let $\varepsilon\in(0,1)$, $\theta\in\Theta$, $n\in\mathbb{N}$ and $N\in\mathbb{N}$. Let $\widehat{\nabla V}(\theta)^{(1)},\ldots,\widehat{\nabla V}(\theta)^{(N)}$ be NN independent copies of the policy gradient estimator defined in (3.6), where each estimator $\widehat{\nabla V}(\theta)^{(k)}$ is computed using an independent version of the estimators $\left(\widehat{\nabla_{\theta}l_{t}^{\theta}}\right)_{t=0}^{T}$ defined in (3.4). Let $\widehat{\nabla V}(\theta)=N^{-1}\sum_{k=1}^{N}\widehat{\nabla V}(\theta)^{(k)}$ be the averaged estimator. Then, for all $\theta\in\Theta$,
$\mathrm{MSE}(\theta,N,n,\varepsilon)\coloneq\max_{j=1,\ldots,D}\mathbb{E}\left[\left|\widehat{\nabla V}(\theta)_{j}-\nabla_{\theta}V(\mu,\theta)_{j}\right|^{2}\right]\leq\mathcal{K}_{1,\varepsilon}\varepsilon^{2}+\frac{\mathcal{K}_{2,\varepsilon}}{N\varepsilon^{2}}+\frac{\mathcal{K}_{3,\varepsilon}}{nN\varepsilon^{4}}\sum_{t=0}^{T}\mathcal{C}_{t}\varepsilon^{-2t},$ (3.9)
@@
$\mathcal{K}_{2,\varepsilon}$ $=2(T+1)^{3}M_{0}^{2}\sum_{t=0}^{T}\left(E_{t}(\theta,\varepsilon)+\left\|\nabla_{\theta}l_{t}^{\theta}\right\|+\varepsilon C\right),$
@@ 4.1 Two-State Two-Action Toy Problem
where $W_{1}$ is the 1-Wasserstein distance, $\lambda>0$ is a fixed scalar parameter and $B$ is a Bernoulli distribution with parameter $p$ such that $1-\lambda_{0}\leq p\leq\lambda_{1}$. The terminal reward function is $g(x,\mu)=r(x,\mu)$. Although the original problem in [17] is an infinite-horizon discounted problem, we consider here a finite-horizon version of with time horizon $T=2$. In this setting, there exists an optimal stationary policy given by
@@
In our experiments, we take the following parameters: $\lambda_{0}=0.5,\lambda_{1}=0.8,\lambda=10,p=0.6$. This policy guarantees that, no matter the initial distribution $\mu_{0}$, the distribution at all time steps $t\geq 1$ is $\mu_{t}^{\star}=B$. The training and evaluation are done as follows. For each training episode we simulate the population starting from a random initial distribution $\mu\in\mathcal{P}(\mathcal{X})^{*}$ by drawing $\mu(L)\sim\mathcal{U}([0.1,0.9])$. Every 10 training episodes, we freeze the policy and sample a validation episode for which we compute the population reward $V(\mu_{0})$ starting from a fixed initial distribution $\mu_{0}=(\mu_{0}(0),\mu_{0}(1))=(0.2,0.8)$. During training, we compute $N=200$ trajectories for MF-REINFORCE, with $n=10$ trajectories for the gradient of logits estimation. For this first example, the policy considered is a simple static policy that outputs a $|\mathcal{X}|\times|\mathcal{A}|$ matrix corresponding to the probability of each action given each state. We make this particular choice of policy to check whether MF-REINFORCE is capable of recovering the optimal static policy given in (4.1). We use the Adam optimizer [24] and train for 5,000 episodes using a learning rate of $10^{-3}$. The experiments are run for $\varepsilon$ ranging in the set $\{0.2,0.5,1.0,2.0\}$. The validation rewards for the different values of $\varepsilon$ are shown in Figure 1. One can see that a larger value of $\varepsilon$ leads to faster increase in the value function at first, but yields worse final policies. In particular, we see that choosing $\varepsilon$ too large can lead to catastophic failure. To evaluate how well the learned policies match the optimal static policy $\pi^{\star}$ (4.1), we compute the average absolute errors in the resulting estimates of $\pi(\mathrm{ST}\mid 0)$ and $\pi(\mathrm{ST}\mid 1)$ over 5 independent training runs for each value of $\varepsilon$ considered. The results are reported in Table 1. These results match adequately with the curves of Figure 1: the smaller $\varepsilon$, the better and closer to optimal the learned policy is.
@@ 4.3 Distribution Planning Example
Finally, we consider a problem similar to the distribution planning example of [6], Section 6.2. The goal is to match a fixed target distribution $\mu_{\mathrm{target}}\in\mathcal{P}(\mathcal{X})$, where $\mathcal{X}=\mathbb{Z}/10\mathbb{Z}$. The action space is $\mathcal{A}=\{-1,0,+1\}$, and the transition probability is deterministic $P(\cdot\mid x,a,\mu)=\delta_{x+a}$. At each step, we penalize the deviation of the population from the target distribution and penalize when an agent chooses to move. Specifically, the step reward is $r(x,a,\mu)=-\lambda|a|-\sum_{x\in\mathcal{X}}|\mu(x)-\mu_{\mathrm{target}}(x)|^{2}$, where $\lambda>0$, and the terminal reward is $g(x,\mu)=-\sum_{x\in\mathcal{X}}|\mu(x)-\mu_{\mathrm{target}}(x)|^{2}$. In this example, we take the length of episodes to be $T=5$ (for both training and validation) and $\lambda=0.01$. For each training episode, we simulate the population starting from a random initial distribution $\mu\in\mathcal{P}(\mathcal{X})^{*}$. Every 10 training episodes, we freeze the policy and sample a validation episode, for which we compute the population reward $V(\mu_{0})$ starting from a fixed initial distribution $\mu_{0}=\mathcal{U}(\mathcal{X})$. During training, we compute $N=500$ trajectories for MF-REINFORCE, with $n=10$ trajectories for the gradient of logits estimation. The policy is given by a 2-layer MLP, with 256 hidden units and $\tanh$ activation. We use the Adam optimizer and train for 100,000 episodes using a learning rate of $10^{-4}$. The experiments are run for $\varepsilon$ ranging in the set $\{0.5,0.75,1.0,2.0\}$. In this example, we see that taking a large value of $\varepsilon$, i.e., $\varepsilon\in\{1,2\}$, leads to more stable training and faster convergence, as shown in Figure 4. However, the resulting trajectories still differ qualitatively between $\varepsilon=1.0$ and $\varepsilon=2.0$, as seen in Figure 5.
@@ 5.1 Proofs of Results of Section 2
$\nabla_{\theta}V(\mu_{0},\theta)$ $=\nabla_{\theta}\int\left(\sum_{t=0}^{T-1}r(x_{t},a_{t},\mu_{t}^{\theta})+g(x_{T},\mu_{T}^{\theta})\right)P_{\theta}({\mathrm{d}}x_{0},da_{0},\ldots,{\mathrm{d}}a_{T-1},{\mathrm{d}}x_{T})$ (5.1)
@@
$\quad+\int\left(\sum_{t=0}^{T-1}r(x_{t},a_{t},\mu_{t}^{\theta})+g(x_{T},\mu_{T}^{\theta})\right)\nabla_{\theta}P_{\theta}({\mathrm{d}}x_{0},da_{0},\ldots,{\mathrm{d}}a_{T-1},{\mathrm{d}}x_{T}).$
@@
$\qquad=\int\left(\sum_{t^{\prime}=0}^{T-1}\nabla_{\theta}\log p(\theta,a_{t^{\prime}},t^{\prime},x_{t^{\prime}},\mu_{t^{\prime}}^{\theta})+\left(\nabla_{l}\log\tilde{p}(\theta,a_{t^{\prime}},t^{\prime},x_{t^{\prime}},l_{t^{\prime}}^{\theta})\tilde{P}(x_{t^{\prime}+1}\mid x_{t^{\prime}},a_{t^{\prime}},l_{t^{\prime}}^{\theta})\right)\nabla_{\theta}l_{t^{\prime}}^{\theta}\right)$
$\qquad\quad\cdot\left(\sum_{t=0}^{T-1}r(x_{t},a_{t},\mu_{t}^{\theta})+g(x_{T},\mu_{T}^{\theta})\right)P_{\theta}({\mathrm{d}}x_{0},{\mathrm{d}}a_{0},\ldots,{\mathrm{d}}a_{T-1},{\mathrm{d}}x_{T}),$
@@
$\mathrm{RF}(\theta)\coloneq\int\left(\sum_{t^{\prime}=0}^{T-1}\nabla_{\theta}\log p(\theta,a_{t^{\prime}},t^{\prime},x_{t^{\prime}},\mu_{t^{\prime}}^{\theta})\right)\left(\sum_{t=0}^{T-1}r(x_{t},a_{t},\mu_{t}^{\theta})+g(x_{T},\mu_{T}^{\theta})\right){\mathrm{d}}P_{\theta},$ (5.2)
@@
By performing the change of variables $\tilde{\lambda}_{t}=l^{\pi}_{t}+\varepsilon\lambda_{t}$, the perturbed value function (2.4) can be written as
@@
$\quad\cdot\mu_{0}({\mathrm{d}}y_{0})\prod_{\tau=0}^{T-1}\mathrm{Leb}({\mathrm{d}}\tilde{\lambda}_{\tau})\tilde{\pi}({\mathrm{d}}a_{\tau}\mid\tau,y_{\tau},\tilde{\lambda}_{\tau})\tilde{P}({\mathrm{d}}y_{\tau+1}\mid y_{\tau},a_{\tau},\tilde{\lambda}_{\tau}).$
Now, let’s consider a policy $\pi_{\theta}$ for $\theta\in{\mathbb{R}}^{D}$. We write $V_{\varepsilon}(\mu,\theta)=V_{\varepsilon}^{\pi_{\theta}}(\mu)$, $\mu_{t}^{\theta}=\mu_{t}^{\pi_{\theta}}$, $l_{t}^{\theta}=l_{t}^{\pi_{\theta}}$, $\alpha_{t}^{\theta,\varepsilon}=\alpha_{t}^{\pi_{\theta},\Lambda,\varepsilon}$ $Y_{t}^{\theta,\varepsilon}=Y_{t}^{\pi_{\theta},\Lambda,\varepsilon}$. The differentiability of $V_{\varepsilon}(\mu,\theta)$ follows from the differentiability of $\tilde{p}$ in $\theta$ and $l$ and the differentiability of $\tilde{P}$ in $l$. Then, given the assumptions on $r,g,\pi_{\theta}$, one can differentiate under the integral sign to get
@@
$\quad\cdot\sum_{t=0}^{T}\varepsilon^{-2}(l_{t}^{\theta}-\tilde{\lambda}_{t})\nabla_{\theta}l_{t}^{\theta}\prod_{\tau=0}^{T}\frac{1}{(2\pi)^{(d-1)/2}\varepsilon^{d-1}}\exp\left(-\frac{\|\tilde{\lambda}_{\tau}-l_{\tau}^{\theta}\|^{2}}{2\varepsilon^{2}}\right)$
$\quad\cdot\mu_{0}({\mathrm{d}}y_{0})\prod_{\tau=0}^{T-1}\mathrm{Leb}({\mathrm{d}}\tilde{\lambda}_{\tau})\tilde{\pi}_{\theta}({\mathrm{d}}a_{\tau}\mid\tau,y_{\tau},\tilde{\lambda}_{\tau})\tilde{P}({\mathrm{d}}y_{\tau+1}\mid y_{\tau},a_{\tau},\tilde{\lambda}_{\tau})$
$\quad+\int\left(\sum_{t=0}^{T-1}\tilde{r}(y_{t},a_{t},\tilde{\lambda}_{t})+\tilde{g}(y_{T},\tilde{\lambda}_{T})\right)\prod_{\tau=0}^{T}\frac{1}{(2\pi)^{d/2}\varepsilon^{d}}\exp\left(-\frac{\|\tilde{\lambda}_{\tau}-\mu_{\tau}^{\theta}\|^{2}}{2\varepsilon^{2}}\right)$
$\quad\cdot\nabla_{\theta}\left(\mu_{0}({\mathrm{d}}y_{0})\prod_{\tau=0}^{T-1}\mathrm{Leb}({\mathrm{d}}\tilde{\lambda}_{\tau})\tilde{\pi}_{\theta}({\mathrm{d}}a_{\tau}\mid\tau,y_{\tau},\tilde{\lambda}_{\tau})\tilde{P}({\mathrm{d}}y_{\tau+1}\mid y_{\tau},a_{\tau},\tilde{\lambda}_{\tau})\right).$
@@
$\qquad=\mu({\mathrm{d}}y_{0})\sum_{t=0}^{T-1}\nabla_{\theta}\log\tilde{p}(\theta,a_{t},t,y_{t},\tilde{\lambda}_{t})\prod_{\tau=0}^{T-1}\mathrm{Leb}({\mathrm{d}}\tilde{\lambda}_{\tau})\tilde{\pi}_{\theta}({\mathrm{d}}a_{\tau}\mid\tau,y_{\tau},\tilde{\lambda}_{\tau})\tilde{P}({\mathrm{d}}y_{\tau+1}\mid y_{\tau},a_{\tau},\tilde{\lambda}_{\tau}).$
@@
$\quad\cdot\left(\sum_{t=0}^{T}\varepsilon^{-1}\Lambda_{t}\left(\nabla_{\theta}l_{t}^{\theta}\right)\right)\Bigg]$
$\quad+\mathbb{E}\Bigg[\left(\sum_{t=0}^{T-1}\tilde{r}(Y^{\theta,\varepsilon}_{t},\alpha_{t}^{\theta,\varepsilon},l_{t}^{\theta}+\varepsilon\Lambda_{t})+\tilde{g}(Y_{T}^{\theta,\varepsilon},l_{T}^{\theta}+\varepsilon\Lambda_{T})\right)$
$\quad\cdot\left(\sum_{t=0}^{T-1}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta,\varepsilon},t,Y_{t}^{\theta,\varepsilon},l_{t}^{\theta}+\varepsilon\Lambda_{t}\right)\Bigg].$
@@
Let’s decompose the expression of the gradient of the perturbed value function (2.6),
@@
$\qquad\to\mathbb{E}\Bigg[\left(\sum_{t=0}^{T-1}r(X^{\theta}_{t},\alpha_{t}^{\theta},\mathbb{P}_{X_{t}^{\theta}})+g(X_{T}^{\theta,\varepsilon},\mathbb{P}_{X^{\theta}_{T}})\right)\left(\sum_{t=0}^{T-1}\nabla_{\theta}\log p(\theta,\alpha_{t}^{\theta},t,X_{t}^{\theta},\mathbb{P}_{X^{\theta}_{t}})\right)\Bigg]=\mathrm{RF}(\theta).$
We now look at the first expectation in (5.3) as $\varepsilon\to 0$, which we rewrite using the following lemma.
@@
$M_{\varepsilon}(\theta)$ $\coloneq\mathbb{E}\left[\left(\sum_{t=0}^{T-1}\tilde{r}(Y^{\theta,\varepsilon}_{t},\alpha_{t}^{\theta,\varepsilon},l_{t}^{\theta}+\varepsilon\Lambda_{t})+\tilde{g}(Y_{T}^{\theta,\varepsilon},l_{T}^{\theta}+\varepsilon\Lambda_{T})\right)\cdot\left(\sum_{t=0}^{T}\varepsilon^{-1}\Lambda_{t}\nabla_{\theta}l_{t}^{\theta}\right)\right]$
$=\mathbb{E}\left[\left(\sum_{t=0}^{T-1}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta}+\varepsilon\Lambda_{t})+\tilde{g}(X_{T}^{\theta},l^{\theta}_{T}+\varepsilon\Lambda_{T})\right)\cdot\left(\sum_{t=0}^{T}\varepsilon^{-1}\Lambda_{t}\nabla_{\theta}l_{t}^{\theta}\right)R_{T}^{\theta,\varepsilon}\right],$
@@
$\quad\cdot\left(1+\varepsilon\nabla_{l}\log\tilde{P}(X_{\tau+1}^{\theta}\mid X_{\tau}^{\theta},\alpha_{\tau}^{\theta},l_{\tau}^{\theta})\cdot\Lambda_{\tau}+o(\varepsilon)\right)\Big]$